Math 141-99 Winter 2019 : Determine the x3.2.RA-9 City A 225 125 1.75 000 175 CtyB 1.50 100 150 100 1.50 Round to the nearest cent as needed)
Answer #1

The data for City A is

{2.25,1.25,1.75,0.00,1.75}  

Mean = sum (Xi)/ n

= (2.25+1.25+1.75+0.00+1.75)/5

= 1.4

Variance = sum(Xi- x bar)^2 / n

= {(2.25-1.4)^2 + (1.25-1.4)^2 + (1.75-1.4)^2 + (0.00-1.4)^2 + (1.75-1.4)^2}/5

= (0.7225+ 0.0225 + 0.1225 + 0 + 0.1225) /5

= 0.198

Thus, Variance for A= $ 0.198

Rounding off to $ 0.20 or 20 cents